A gas exist at 3.1 atm of pressure with a volume of 36.5 L and a temperature of 30°C. What is the volume of this gas in liters when it is placed in an environment under STP?


1
Expert's answer
2021-03-29T05:15:51-0400

P = 3.1 atm , V = 36.5L , R = 0.0821 , T = 30⁰C = 303K

n = PV/RT

n = 3.1×36.5 / 0.0821×303

n = 4.54moles

At STP 1 mole occupies 22.4 L

4.54 moles occupies = 22.4×4.54

= 101.7L